OpenStudy (anonymous):

CAN ANYONE SOLVE? Evaluate: 6 + (-3)^2 + 6 * 5^2/ 6^2 * (6-5)

OpenStudy (anonymous):

just.. solve it? i can check the answer for you

OpenStudy (anonymous):

ok i got 180/35

OpenStudy (anonymous):

\[6 + (-3)^2 + 6 \times \frac{5^2}{6^2} \times (6-5)\] right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this one?\[\frac{6+ (-3)^2 + 6 \times 5^2}{6^2 \times (6-5)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then your answer is wrong

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i just simplified it\[\frac{6+9 + 6 \times 25}{36 \times 1}\]

OpenStudy (anonymous):

whats your final answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{15+150}{36}=\frac{165}{36}=\frac{55}{12}\]

OpenStudy (anonymous):

thanks so much!